OpenStudy (anonymous):

The circulation of a newsletter decreased from 3,200 to 2,464. What was the percent of decrease in circulation?

OpenStudy (anonymous):

i think its 23.

OpenStudy (anonymous):

23%* :P

OpenStudy (anonymous):

2464/3200, so 1-.77=0.23 23%

OpenStudy (anonymous):

\[\frac{3200-2464}{3200}=0.23 = 23\% \text{ decrease} \]

OpenStudy (anonymous):

thanks everyone :)

OpenStudy (anonymous):

yay i got it right!!